OpenStudy (anonymous):

What is the greatest common divisor of 108, 72, and 48?

OpenStudy (freckles):

well lets look at gcd(108,72) do you know how to write the prime factorization for 108 and 72?

OpenStudy (freckles):

or do you want to use euclidean's algoritm

OpenStudy (anonymous):

What is the second one and the first one?!?!?!?!? im only in sixth grade EVERYTHING is confusing

OpenStudy (freckles):

well we can write the prime factorization for 108 and 72

OpenStudy (freckles):

euclidean's algorithm is normally thought in advanced math classes

OpenStudy (freckles):

taught*

OpenStudy (freckles):

you know what a prime number is?

OpenStudy (anonymous):

yes

OpenStudy (freckles):

ok so can you write the prime factorization for 108?

OpenStudy (anonymous):

um just a sec let me think

OpenStudy (anonymous):

2^2.3^3?!?!?!?!?!?

OpenStudy (freckles):

good job let's just compare the first two numbers for right now 108=2(54)=2(6)(9)=2(2)(3)(3)(3) 72=8(9)=(2)(2)(2)(3)(3)

OpenStudy (freckles):

how many pairs of 2's do they have in common?

OpenStudy (anonymous):

one

OpenStudy (anonymous):

i think

OpenStudy (freckles):